### v1

# F-LMM: Grounding Frozen Large Multimodal Models

Abstract
Endowing Large Multimodal Models (LMMs) with visual grounding capability can significantly enhance AIs’ understanding of the visual world and their interaction with humans. However, existing methods typically fine-tune the parameters of LMMs to learn additional segmentation tokens and overfit grounding and segmentation datasets. Such a design would inevitably cause a catastrophic diminution in the indispensable conversational capability of general AI assistants. In this paper, we comprehensively evaluate state-of-the-art grounding LMMs across a suite of multimodal question-answering benchmarks, observing pronounced performance drops that indicate vanishing general knowledge comprehension and weakened instruction following ability. To address this issue, we present F-LMM—grounding frozen off-the-shelf LMMs in human-AI conversations—a straightforward yet effective design based on the fact that word-pixel correspondences conducive to visual grounding inherently exist in the attention weights of well-trained LMMs. Using only a few trainable CNN layers, we can translate word-pixel attention weights to mask logits, which a SAM-based mask refiner can further optimise. Our F-LMM neither learns special segmentation tokens nor utilises high-quality grounded instruction-tuning data, but achieves competitive performance on referring expression segmentation and panoptic narrative grounding benchmarks while completely preserving LMMs’ original conversational ability. Additionally, with instruction-following ability preserved and grounding ability obtained, our F-LMM can perform visual chain-of-thought reasoning and better resist object hallucinations. Code and models will be released at https://github.com/wusize/F-LMM.
Figure 1: An example of user-AI conversation around an image. Left: The current state-of-the-art grounding model GLaMM [57] is effective for grounded conversation when prompted by "answer with interleaved masks", but fails to follow user instruction to answer a single word (yes or no) and misunderstands the question as a referring segmentation prompt. Right: Our F-LMM preserves instruction-following ability while being able to perform visual grounding.
1 Introduction
As one of the pivotal milestones in AGI, recent Large Multimodal Models (LMMs)—integrating Large Language Models (LLMs) with visual signals—have demonstrated remarkable success in multimodal understanding, reasoning and interaction [39, 37, 38, 42, 32, 62, 75]. To further advance LMMs with better perception capability, a recent line of research [77, 28, 57, 58, 67, 80] that visually grounds language contents in user-model conversations has drawn increasing attention. This explicit association between key phrases or words with visual objects greatly enhances LMMs’ understanding of the visual world and allows for more intuitive and meaningful human-AI interactions.
Figure 2: (a) Geometric and spatial cues conducive to visual grounding are observed in the visualisations of word-image attention maps in frozen LMMs. (b) Existing grounding LMMs are fine-tuned to generate a special mask token (e.g., [SEG]) for visual grounding purposes, which ruins the original conversational ability. (c) Our F-LMM translates word-image attention maps from frozen LMMs to grounding masks, while fully preserving the general-purpose chat capability.
By design, one commonly adopted build (Figure 2(b)) for visually grounding language contents is connecting LMMs with a mask head (e.g., Segment Anything Model (SAM) [26]), wherein both the LLM backbone and the mask head are fine-tuned with well-prepared visual grounding data that contains segmentation annotations. Also, some additional learnable tokens (e.g., [SEG]) are introduced to the LMMs’ vocabulary, to directly associate key phrases or words with visual objects in conversations. However, this design will inevitably provoke a catastrophic diminution in general knowledge comprehension and instruction-following ability due to the following reasons. First, existing segmentation and visual grounding data only contain elementary patterns for answering simple grounding prompts. Second, during the fine-tuning stage, the LMMs are mainly optimised for effectively modelling the relationship between key phrases or words and special segmentation tokens, i.e., overfitting the segmentation and grounding data. Therefore, the conversational ability that is indispensable in building general AI assistants is sacrificed. For instance, the state-of-the-art grounding model GLaMM [57] fails to answer a simple yes-or-no question (Figure 1). Moreover, quantitative evaluations of existing grounding LMMs in conversational ability are presented in Table 1, with zero or near-zero scores on general multimodal question-answering benchmarks necessitating instruction-following ability.
To deal with this dilemma, one possible option is to collect high-quality training data encompassing both meaningful conversations and mask annotations. For example, LLaVA-G [77] annotates the 150k LLaVA-Instruct data samples [39] with segmentation masks so that the LMMs simultaneously learn to chat and segment. Nonetheless, annotating high-quality grounded conversation data is costly and hard to scale up. Despite being trained on the costly annotated data, LLaVA-G still lags behind general-purpose LMMs on multimodal understanding tasks. Furthermore, training on large-scale annotated data normally consumes significant computational resources, which is, obviously not a resource-efficient solution.
In this paper, we propose a straightforward yet effective design, i.e., grounding frozen LMMs (dubbed as F-LMM) in human-AI conversations. Thinking from first principles, we argue that freezing the parameters of well-trained LMMs is the most feasible design choice for completely maintaining the original excellent conversational ability when building general-purpose grounding LMMs. In particular, we take inspiration from the built-in interpretability of the attention mechanism in transformers [65, 6] that represents interrelations between text and image contents in design. We observe that off-the-shelf LMMs already produce word-pixel correspondences necessary for visual grounding, despite not being directly pre-trained with region or pixel annotations. As exhibited in Figure 2(a), we visualise word-image attention maps from frozen LMMs via K-Means clustering, demonstrating prominent geometric and spatial cues of the objects.  [^1] For better visibility, we perform K-Means clustering on the stack-up of all attention maps collected in a forward pass instead of selecting a single attention map. For example, coarse visual grounding masks for key phrases (e.g., "two girls", "two elephants", and "the plate") in language sentences can emerge from attention maps in LMMs. Therefore, our F-LMM takes these visual-language correspondences as useful segmentation priors for decoding grounding masks, neither further tuning the LMMs’ weights nor learning a special segmentation token to model object locations, as shown in Figure 2(c).
Notably, the only trainable part of our F-LMM is a mask head. It comprises a CNN-based mask decoder (a tiny U-Net [59]) that translates stacked attention maps to mask logits and a light-weight mask refiner (retrofitted from SAM [26]’s mask head) that uses additional image and language cues to refine the semantic-agnostic masks from the mask decoder. Moreover, we only use the RefCOCO(+/g) [23, 44] and PNG [17] datasets as our training data, enabling LMMs to segment user-described objects and ground key phrases or words in a text sequence. Unlike previous works [77, 57, 58], our F-LMM eliminates the necessity for high-quality conversation data that are annotated with masks to preserve conversational ability when learning grounding.
In experiments, our F-LMM maintains the original excellence of off-the-shelf LMMs on general question-answering benchmarks, while achieving competitive results on referring segmentation and phrase grounding. Compared with existing grounding LMMs, F-LMM offers the best balance between grounding and chat capabilities. In addition, with instruction-following ability preserved and grounding ability obtained, F-LMM unleashes visual chain-of-thought reasoning in a zero-shot manner by exhibiting improvement on the VisCoT benchmark [60] and better resists object hallucinations on the POPE benchmark [34].
2 Related Work
Large Multimodal Models. Recent advancements in LMMs  [2, 31, 14, 39, 37, 38, 72, 3, 30, 35, 45, 42, 32, 62, 29] have been fueled by the success of LLMs [4, 1, 78, 63, 64, 22, 12, 49, 46] since the debut of GPT series [54, 55, 4, 1] that feature an auto-regressive framework based on transformer decoder [65]. These LLMs possess general world knowledge and excellent conversational ability to follow human instructions, thanks to large-scale generative pre-training [4] and supervised finetuning on instruction-tuning data [68] or human feedback [48]. By integrating image representations from vision encoders [53, 76] to LLMs, LMMs enable visual understanding and reasoning as AI assistants. This integration is usually established by a multilayer perceptron (MLP) that directly maps image features to the LLMs’ input embedding space [39, 37, 38, 42, 32, 62, 75] or a cross-attention module that abstracts the image contents with a set of query embeddings [2, 31, 3, 72]. In our research, we build F-LMM on LMMs of the former type (MLP-based), which preserves images’ 2-D topological structure in the cross-modal integration.
Visual Segmentation. The task of predicting 2D masks for visual objects is known as image segmentation, which can be categorised into semantic segmentation [8, 5, 81, 11], instance segmentation [20, 10, 79] and panoptic segmentation [25, 9, 70, 33] depending on whether the goal is to differentiate pixel semantics or object instances. These standard segmentation approaches rely on a pre-defined set of object classes for recognition. In contrast, referring expression segmentation (RES) [23, 44, 47, 82, 43, 71, 36] involves segmenting objects based on free-form human language descriptions, allowing for enhanced human-model interaction. Additionally, panoptic narrative grounding (PNG) [17, 15, 66, 18] requires segmenting masks for key phrases or words in a sentence. In this study, we leverage RES and PNG tasks to evaluate the grounding capability of LMMs. In addition, the prompt-based SAM [26] pre-trained on billion-scale high-quality mask data have become a constituent component in many grounding LMMs to boost segmentation performance. We also adopt SAM’s mask head to initialise our mask refiner.
Grounding Large Multimodal Models. Grounding Large Multimodal Models [50, 7, 3, 73, 77, 80, 58, 57, 67, 28, 69, 51, 75] can localise language contents during user-model conversations. Some approaches [50, 7, 73, 3] represent coordinates of bounding boxes as texts and train LMMs to predict the coordinates in a generative manner. Several recent works [28, 77, 58, 67, 57, 51] train LMMs to predict a special segmentation token for encoding the grounded object and utilise a segmentation head (e.g., SAM [26]) to decode object masks. This study mainly focuses on grounding LLMs with segmentation ability for visual perception. To obtain competitive visual grounding performance, existing works extensively fine-tune the parameters of LMMs on a large amount of segmentation [81, 5, 25, 56, 19] and grounding [23, 44, 24, 27, 52, 17] datasets. And to balance the LMMs’ grounding and conversational abilities, there are efforts [77, 58, 80] to collect high-quality instruction-tuning data annotated with segmentation masks. In contrast, we make the first attempt to build grounding LMMs on top of off-the-shelf LMMs without fine-tuning their parameters. Furthermore, we bypass the need for grounded instruction-tuning data to preserve decent chat ability.
3 Method
Figure 3: The overall pipeline of F-LMM. The word-image attention maps from the frozen LLM serve as segmentation priors for the mask head. The mask head encompasses a mask decoder that translates attention weights to mask logits and a mask refiner that optimises the mask decoder’s predictions. $M$ and $N$ represent the numbers of transformer layers and attention heads.
In this section, we introduce our F-LMM by first probing the causal attention mechanism in LMMs with visualisations of word-image attention maps in Sec 3.1. Then, we elaborate on F-LMM exploiting segmentation priors from frozen LMMs for visual grounding using the mask head in Sec 3.2. Finally, we show how to perform referring expression segmentation and phrase grounding with our F-LMM for human-AI conversations in Sec 3.3. The overall pipeline is illustrated in Figure 3.
3.1 Segmentation Priors from Frozen LMM
Vision-Language Sequence. A typical build of a Large Multimodal Model (LMM)  [^2] In this paper, the term ‘multimodal’ stands for vision and language modalities. comprises an image encoder $f_{v}$ (e.g., CLIP [53]  [^3] The image encoder might be any vision model that is pre-trained on image-text pairs. We use the classic term ‘CLIP’ in this paper to represent all such models for brevity.), a vision-language projector $f_{p}$, and a Large Language Model (LLM) $f_{\mathrm{llm}}$. The inputs to an LMM are usually an image ${\mathbf{X}}_{v}\in{\mathbb{R}}^{3\times H\times W}$ and the associated text ${\mathbf{X}}_{t}$. The input image is first encoded by the vision encoder $f_{v}$ and then mapped to the input space of the LLM $f_{\mathrm{llm}}$ by the projector $f_{p}$:
${\mathbf{Z}}_{v}=f_{p}(\texttt{Flatten}(f_{v}({\mathbf{X}}_{v})))\in{\mathbb{R}}^{hw\times d},$ (1)
where $h$ and $w$ are the height and width of projected feature maps via $f_{v}$. The Flatten operation unfolds the 2-D image feature map to a 1-D sequence. The constant $d$ is the hidden state dimension of the LLM $f_{\mathrm{llm}}$. Likewise, the text input is first encoded as discrete tokens and then mapped to text embeddings:
${\mathbf{Z}}_{t}=\texttt{Embed}(\texttt{Tokenize}({\mathbf{X}}_{t}))\in{\mathbb{R}}^{L\times d},$ (2)
where $L$ denotes the length of text embeddings. The visual-language sequence input to the LLM $f_{\mathrm{llm}}$ is a concatenation of image and text embeddings: ${\mathbf{Z}}=\{{\mathbf{Z}}_{v},{\mathbf{Z}}_{t}\}\in{\mathbb{R}}^{(hw+L)\times d}$.
Segmentation Priors in Self-Attention. The vision-language sequence is mainly processed by causal self-attentions [65, 54] in the LLM $f_{\mathrm{llm}}$, including inner product and weighted-sum operations. Specifically, for a word token with position index $i$ in the vision-language sequence ${\mathbf{Z}}$ , its embedding ${\bm{z}}^{i}$ is updated by the weighted sum of the first $i$ embeddings: $\hat{{\bm{z}}}^{i}=\texttt{SoftMax}(\frac{{\bm{z}}^{i}\cdot{\mathbf{Z}}[:i]}{d})\cdot{\mathbf{Z}}[:i]$, where $\texttt{SoftMax}(\frac{{\bm{z}}^{i}\cdot{\mathbf{Z}}[:i]}{d})$ is the attention weights. Here, we omit the residual layers and feedforward layers for brevity.
Figure 4: Visualisations of word-image attention maps. The letters $m$ and $n$ indicate that the attention map is derived from the $n$-th attention head of the $m$-th transformer layer. Though noisy, the objects are observable in the attention maps. The visibility is further enhanced when we stack up all the attention maps and perform KMeans clustering.
Considering the word-image interaction in the multimodal scenario, we can select the word token’s attention weights with the image embeddings from the overall vision-language attention weights:
$\bm{a}^{i}=\texttt{Unflatten}(\texttt{SoftMax}(\frac{{\bm{z}}^{i}\cdot{\mathbf{Z}}[:i]}{d})[:hw])\in{\mathbb{R}}^{h\times w},$ (3)
where Unflatten restores the 2-D spatial structure from the 1-D sequence to form an attention map. In Figure 4, we visualise such word-image attention maps from various transformers layers and attention heads in an LMM (i.e., DeepseekVL-1.3B [42]). The objects’ shape and location can be observed in word-image attention maps of certain layers or heads. The visibility is further enhanced when we stack the attention maps from all layers and heads and perform K-Means clustering. It can be observed that the attention maps offer meaningful segmentation priors with spatial and geometric cues for grounding objects visually.
Language Cues. In addition to the spatial and geometric cues from word-image attention maps, F-LMM can also capitalise on the object’s corresponding text embeddings from the LLM $f_{\mathrm{llm}}$, which provide extra language cues for the grounding of visual objects.
3.2 Visual Grounding with Mask Head
We use the segmentation priors from the frozen LMM for pixel-level grounding, with the help of a mask head consisting of a mask decoder and a mask refiner.
Mask Decoder. The mask decoder $f_{d}$ is a 2-D CNN model that transforms the word-image attention maps of grounded objects into mask logits, which is instantiated by a 3-stage U-Net [59]. Please refer to Sec A.2 of the appendix for details of the mask decoder. The extraction of word-image attention map $\bm{a}^{i}$ for a word token with position index $i$ is illustrated in Eq. 3 and Figure 3. For an object described by multiple words, we merge their corresponding word-image attention maps to a single attention map $\bm{a}$ via element-wise average or max operation. The attention map $\bm{a}$ is further normalised as $\bm{a}/\mathrm{sum}(\bm{a})$ so that all elements sum to 1. Considering $M$ layers and $N$ attention heads, we stack the $MN$ attention maps as $\bm{A}\in{\mathbb{R}}^{MN\times h\times w}$, which forms the input to a mask decoder. Given the importance of high input resolution for segmentation models, we upsample the stacked attention maps $\bm{A}$ to $h^{\prime}\times w^{\prime}$ by bilinear interpolation before feeding it to a mask decoder, where $h^{\prime}>h$ and $w^{\prime}>w$. In practice, we set $h^{\prime}=w^{\prime}=64$. Then, the mask decoder maps $\bm{A}$ into mask logits: ${\mathbf{M}}_{\mathrm{logits}}=f_{d}(\bm{A})$. We derive the corresponding binary mask via ${\mathbf{M}}_{\mathrm{pred}}={\mathbf{M}}_{\mathrm{logits}}>0$. During training, the mask decoder is optimised with BCE and DICE losses [61].
Mask Refiner. The mask refiner $f_{r}$ is retrofitted from the mask head of SAM [26], which predicts masks based on prompts as well as image embeddings from SAM’s ViT-based image encoder. To refine the output of the mask decoder $f_{d}$, we re-use SAM’s prompt encoder to transform ${\mathbf{M}}_{\mathrm{logits}}$ into dense prompt embeddings (i.e., a 2-D feature map) and the bounding box of ${\mathbf{M}}_{\mathrm{pred}}$ to box embeddings. In addition to the location cues from the mask and the box, the language cues, i.e., the object’s corresponding text embeddings, are also utilised by $f_{r}$. Considering text embeddings from the $M$ transformer layers, we train $M$ learnable scalars to calculate a weighted sum of these text embeddings. The weighted-summed text embeddings are processed by a linear layer and then concatenated with the box embeddings to form sparse prompt embeddings. The dense and sparse prompt embeddings, together with SAM’s image embeddings, are fed to the mask refiner $f_{r}$ for finer-grained mask predictions ${\mathbf{M}}^{\prime}_{\mathrm{pred}}$. During training, we keep the ViT-based image encoder of SAM frozen and optimise the mask refiner $f_{r}$ using BCE loss and DICE loss [61]. For more details on the SAM’s prompt-based mask head, please refer to the original SAM paper [26].
3.3 User-AI Interaction with Grounding
We elaborate on how F-LMM works for user-AI conversations in two typical scenarios, i.e., referring expression segmentation and phrase grounding.
Referring Expression Segmentation. In this scenario, the model is supposed to segment user-described objects. Existing works [28, 57] request the LMM generate a special segmentation token in the answer, which is then decoded as the mask of the described object. In our F-LMM, we can directly perform grounding for user descriptions using word-image attention maps and text embeddings.
Phrase Grounding. In user-model conversations, the grounding LMMs can localise key phrases or words when chatting with the user. Unlike existing works [58, 57] that generate special tokens to indicate grounded objects, we use the spaCy toolkit [21] to obtain object words and phrases in the texts and decode grounding mask from frozen LMMs. This disentanglement of text generation and phrase selection in our design also allows users to decide which words or phrases to ground.
4 Experiments
Table 1: The main evaluation results on question-answering benchmarks, referring expression segmentation (RES) benchmark and panoptic narrative grounding (PNG) benchmark. MMB: MMBench; LLaVAW: LLaVA-In-the-Wild; RefC(+/g): RefCOCO(+/g). LLaVA-1.6 and MGM-HD take high-resolution image inputs. LLaVA-1.6-M-7B means the model is based on Mistral-7B [22]. GLaMM-FS-7B means we use the ‘FullScope’ version of GLaMM.
Model Multimodal Question Answering RES PNG
MME MMB MMVet LLaVAW RefC RefC+ RefCg All Thing Stuff
Specialised Segmentation Models
MCN [43] - - - - 62.4 50.6 49.2 54.2 48.6 61.4
LAVT [71] - - - - 72.7 62.1 61.2 - - -
GRES [36] - - - - 73.8 66.0 65.0 - - -
X-Decoder [82] - - - - - - 64.6 - - -
SEEM [83] - - - - - - 65.7 - - -
PNG [17] - - - - - - - 55.4 56.2 54.3
PPMN [15] - - - - - - - 59.4 57.2 62.5
XPNG [18] - - - - - - - 63.3 61.1 66.2
Existing Grounding LMMs
PixelLM-7B [58] 309/135 17.4 15.9 46.4 73.0 66.3 69.3 43.1 41.0 47.9
LISA-7B [28] 1/1 0.4 19.1 47.5 74.9 65.1 67.9 - - -
PerceptionGPT-7B [51] - - - - 75.1 68.5 70.3 - - -
LLaVA-G-7B [77] - - - 55.8 77.1 68.8 71.5 - - -
GroundHog-7B [80] - - - - 78.5 70.5 74.1 66.8 65.0 69.4
GLaMM-FS-7B [57] 14/9 36.8 10.3 32.0 78.6 70.5 74.8 55.8 52.9 62.3
LaSagnA-7B [67] 0/0 0.0 16.7 34.5 76.8 66.4 70.6 - - -
Grounding Frozen General-Purpose LMMs by F-LMM (Ours)
DeepseekVL-1.3B [42] 1307/225 64.6 34.8 51.1 75.0 62.8 68.2 64.9 63.4 68.3
MGM-2B [32] 1341/312 59.8 31.1 65.9 75.0 63.7 67.3 65.6 64.4 68.4
LLaVA-1.5-7B [37] 1511/348 64.3 30.5 69.0 75.2 63.7 67.1 64.8 63.4 68.2
HPT-Air-6B [62] 1010/ 258 69.8 31.3 59.2 74.3 64.0 67.5 65.5 64.0 68.8
HPT-Air-1.5-8B [62] 1476/308 75.2 36.3 62.1 76.3 64.5 68.5 65.4 64.1 68.5
MGM-7B [32] 1523/316 69.3 40.8 75.8 75.7 64.8 68.3 66.3 65.3 68.6
DeepseekVL-7B [42] 1468/298 73.2 41.5 77.8 76.1 66.4 70.1 65.7 64.5 68.5
LLaVA-1.6-7B [38] 1519/322 68.1 44.1 72.3 75.8 65.8 70.1 66.3 65.1 69.0
LLaVA-1.6-M-7B [38] 1501/324 69.5 47.8 71.7 75.7 66.5 70.1 66.5 65.4 69.1
MGM-HD-7B [32] 1546/319 65.8 41.3 74.0 76.1 65.2 68.5 66.7 65.6 69.1
4.1 Implementation Details
Model Architectures. We build F-LMM on several open-sourced LMMs, including LLaVA-1.5 [37], LLaVA-Next [38], MiniGemini [32], DeepseekVL [42] and HPT-Air [62]. The main experiment covers 10 LMMs with model sizes ranging from 1.3B to 8B. We employ a lightweight 3-stage U-Net [59] as the CNN-based mask decoder to transform segmentation priors from frozen LLMs. The U-Net architecture features an encoder-decoder structure with skip connections, wherein feature maps are downsampled in the encoder and upsampled in the decoder. Please check Sec A.2 of the appendix for more details on the mask decoder. As for the SAM-based mask refiner, we choose SAM ViT-L [26] that balances cost and performance well.
Model Training. F-LMM is implemented on XTuner [13]. We train F-LMM on RefCOCO(+/g) [23, 44] and PNG [17] datasets with about 190k data samples on a single machine with 8 NVIDIA A800-40G GPUs, which costs about 20 hours for each round of model training. We set the batch size to 8 and train models for 8 epochs, with gradient clipping at a max norm of 1.0. The AdamW [41] optimiser is used with a learning rate of 1e-4, a weight decay of 0.01, and betas as (0.9, 0.999). We choose a warm-up ratio of 0.03 at the beginning of training to stabilise model optimisation.
4.2 Main Evaluation
Our main evaluations cover both the conversational and grounding ability of LMMs. We summarise the evaluation results of grounding LMMs in Table 1. Please refer to Sec A.1 in the Appendix for more detailed results.
Benchmarks. For comprehensive conversational ability evaluation, we choose four widely used general question-answering benchmarks including MME [16], MMBench [40], LLaVA-In-the-Wild [39] and MMVet [74]. The MME and MMBench require an LMM to strictly follow the instruction to reply with single words (yes or no) or answer MCQs with alphabetical letters (i.e., answering A, B, C, or D). The LLaVA-In-the-Wild and MMVet benchmarks ask a model to respond with open-ended texts while demanding general world knowledge comprehension. In terms of grounding ability evaluation, we assess the LMMs’ ability to segment user-described objects on referring expression segmentation (RES) [23, 44] benchmarks including RefCOCO, RefCOCO+, and RefCOCOg, using the cIoU metric. Due to limited space, we only report results on the Val splits of RefCOCO(+/g) in Table 1. We also test the LMMs’ ability to ground key phrases or words in user-model conversations on the Panoptic Narrative Grounding (PNG) [17] benchmark, measuring individual mask recalls on thing/stuff objects and overall recall scores.
Comparisons with Existing Methods. We compare F-LMM with existing grounding LMMs. As shown in Table 1, our F-LMM provides the best balance with conversational and grounding abilities among compared methods. On the question-answering benchmarks, existing grounding LMMs obtain zero or near-zero scores on MMBench and MME while lagging significantly behind general-purpose LMMs on MMVet and LLaVA-In-the-Wild benchmarks, indicating compromised instruction-following ability and weakened general knowledge comprehension. On the RES and PNG benchmarks, our F-LMM achieves comparable results despite not having the parameters of LMMs fine-tuned for grounding purposes. Compared with standard segmentation models, F-LMM outperforms all these specially designed models on both RES and PNG benchmarks.
Table 2: Unleashing visual chain-of-thought reasoning with both excellent grounding and instruction-following ability.
Model Visual VisCoT Benchmark POPE
CoT DocVQA TextCaps TextVQA DUDE SROIE Infographics Acc F1
VisCoT-7B [60] 47.6 67.5 77.5 38.6 47.0 32.4 86.5 -
F-LMM (Ours)
DeepseekVL-1.3B [42] 30.4 58.2 69.7 23.9 20.0 31.0 87.4 86.6
DeepseekVL-1.3B [42] 38.6 62.2 75.0 31.8 31.6 34.4 88.3 88.1
DeepseekVL-7B [42] 43.2 63.5 74.5 32.0 28.4 43.2 87.0 86.0
DeepseekVL-7B [42] 53.8 67.9 78.4 42.3 44.1 49.1 88.0 87.7
4.3 Unleashing Visual Chain-of-Thought Reasoning
Considering our F-LMM can ground objects without losing any instruction-following ability, we further study whether visual chain-of-thought (CoT) reasoning can be elicited in our model. In human-AI conversations that involve Visual CoT, an LMM first localises the region/object relevant to the human’s question and then generates the final answer by zooming in on the question-related region. Here, we use DeepseekVL-1.3B and DeepseekVL-7B models that support multi-visual inputs and evaluate on the VisCoT benchmark [60]. As shown in Table 2, our models achieve remarkable performance gains when prompted in a visual CoT manner. It is noticeable that our F-LMM even outperforms VisCoT-7B [60] that has been well-tuned on the training set of VisCoT data [60]. Furthermore, we perform visual CoT reasoning on the object hallucination benchmark POPE [34] and observe significant performance gain in resisting object hallucinations. Thanks to the combination of excellent grounding and instruction-following abilities, our F-LMMs have the potential to perform complex visual perception and reasoning. For more details of visual CoT reasoning on the VisCoT benchmark [60], please refer to Figure 6.
Table 3: Ablation study of F-LMM (based on DeepseekVL-1.3B) on the PNG benchmark. The blue colour indicates our default design choice.
((a)) Plain CNN v.s. U-Net
# Mask PNG
Decoder All Thing Stuff
1 Plain CNN 64.5 63.1 67.8
2 U-Net 64.9 63.4 68.3
((b)) Merge Type
# Merge PNG
Type All Thing Stuff
1 Max 64.6 63.1 68.0
2 Average 64.9 63.4 68.3
((c)) Input Normalisation
# Normalize PNG
Input All Thing Stuff
1 64.5 63.0 68.0
2 64.9 63.4 68.3
((d)) Input Size
# Input PNG
Size All Thing Stuff
1 32 64.2 62.8 67.6
2 64 64.9 63.4 68.3
3 128 65.0 63.5 68.4
((e)) SAM Variants
# SAM PNG
Variant All Thing Stuff
1 ViT-B 63.0 61.4 66.8
2 ViT-L 64.9 63.4 68.3
3 ViT-H 65.0 63.5 68.3
((f)) Prompts for Mask Refiner
# Prompts PNG
mask box text All Thing Stuff
1 63.4 62.0 69.8
2 63.7 62.2 67.1
3 64.9 63.4 68.3
Table 4: Analysis of original visual grounding ability in LMMs by discarding object-centred data samples during training.
# Model VisualGenome RES PNG
data RefCOCO RefCOCO+ RefCOCOg All Thing Stuff
1 LLaVA-1.5-7B [37] 75.2 63.7 67.1 64.8 63.4 68.2
2 LLaVA-1.5-7B [37] 73.2 60.8 65.1 64.5 63.0 67.9
4.4 Ablation Study
We investigate the effects of design choices of F-LMM. All the ablation studies are conducted on the PNG benchmark (‘All’) using the smallest LMM, DeepseekVL-1.3B [42].
Mask Decoder. We consider two architecture variants of the mask decoder: a U-Net [59] involves several downsampling and upsampling operations and a plain CNN without variations in the resolution of feature maps. To ensure a fair comparison, we keep both two architectures having approximately the same number of parameters (i.e., 8M). As shown in Table 3, the U-Net outperforms the plain CNN by a 0.4 margin, substantiating the importance of multi-scale structure in building segmentation models. Then, we study the effectiveness of normalizing word-image attention maps, which ensures all attention scores sum up to 1.0. By eliminating the influence of varying sequence length on the magnitude of attention scores, the normalisation of the attention map can provide a performance increase of 0.4, as shown in Table 3. In addition, we also find that averaging attention maps of a multiple-word object achieves better performance than max operation by a margin of 0.3. Finally, we study the input size of the mask decoder in Table 3, including $32\times 32$, $64\times 64$, and $128\times 128$, observing the performance increases with the resolution of inputs. By default, we choose $64\times 64$ to balance cost and performance.
Mask Refiner. We consider three SAM [26] mode variants, i.e., ViT-B(ase), ViT-L(arge) and ViT-H(uge) as mask refiners. As shown in Table 3, the performance grows with model sizes. For a good trade-off between cost and performance, we select ViT-L as the default mask refiner. To analyze the effect of different prompts on the mask refiner, we start by only using coarse mask logits from the CNN-based mask decoder. Then, we gradually add the bounding box of the coarse masks and text embeddings as prompts. As shown in Table 3, the performance increases when we stack up more prompts for mask refinement. It is evident that utilizing the coarse mask logits only can already yield considerable performances, verifying that the geometric and spatial cues from the transformers’ attention mechanism are effective for visual grounding.
4.5 Analysis and Visualisation
Model Chat Ground
DeepseekVL-1.3B [42] 7.75 8.33
MGM-2B [32] 6.00 8.33
LLaVA-1.5-7B [37] 6.75 7.83
HPT-Air-6B [62] 9.00 7.16
HPT-Air-1.5-8B [62] 6.50 7.00
MGM-7B [32] 5.75 4.83
DeepseekVL-7B [42] 3.75 4.00
LLaVA-1.6-7B [38] 2.75 3.00
LLaVA-1.6-M-7B [38] 3.25 1.66
MGM-HD-7B [32] 3.50 2.83
Figure 5: The left table shows the average ranks of each LMM on question-answering (‘Chat’) and grounding benchmarks (‘Ground’). The dashed line in the right figure is the linear fit of the rank data points, indicating a positive correlation between abilities to chat and ground.
Figure 6: Examples of visual chain-of-thought reasoning. The model used is DeepseekVL-1.3B and the samples are taken from the test set of VisCoT data [60]. The LMM is first prompted to think about the question-related object, which is grounded by the mask head of F-LMM. The region of the question-related object is cropped and fed to the LMM to help answer the question.
The Origin of Visual Grounding Ability. One might question whether the visual grounding ability in LMMs originates from object-centred training samples. For example, a considerable part of the LLaVA-Instruct data [37] is from VisualGenome [27] dataset in which question-answer pairs are based on specific regions or objects. Regarding this concern, we discard the data samples that are taken from the VisualGenome and train an LLaVA-1.5-7B model on the rest of the LLaVA-Instruct data. As shown in Table 4(#2), we can still obtain competitive performance on the grounding benchmarks. Therefore, we believe the built-in grounding ability in LLMs’ attention mechanism can be learned even if the LMMs are trained only on image-level data samples.
Better Chatting Means Better Grounding? We study the correlation between performance on the question-answering benchmarks and the grounding benchmarks. For the 10 models reported in Table 1, we calculate their average ranks in each benchmark category (the lower the rank, the better the performance). In Figure 5, we visualise the two types of ranks as 2D coordinates, i.e., (Chat Rank, Ground Rank), and linearly fit these rank data points. As indicated by the blue dashed line, frozen LMMs with better conversational ability would serve as better backbones for grounding. We also observe that larger LMMs tend to be better at both conversation and grounding, and LMMs with larger input resolution (e.g., LLaVA-1.6 and MGM-HD) can handle both tasks better.
Visualisation of Visual CoT. In Figure 6, we show examples of visual chain-of-thought reasoning (Visual CoT), in which the LMM is first prompted to answer "What object is the most relevant to the question?". Then, the mask head of F-LMM grounds the LMM’s answer about the relevant object by generating a segmentation mask, the bounding box of which is used to crop the object region from the original image. Finally, the cropped image region is fed into the LMM to obtain the final answer. As shown in Figure 6, the Visual CoT powered by the LMM’s grounding ability is helpful when the LMM needs to focus on the question-related regions for better visual perception and reasoning.
5 Conclusion
In this work, we studied the limitation of existing grounding LMMs, i.e., the loss of general world knowledge and instruction-following ability in the course of seeking state-of-the-art performance on grounding tasks. Regarding this issue, we make the first attempt to ground completely frozen LMMs that are well-trained for user-model conversation, based on the observation that geometric and spatial cues necessary for visual grounding already exist in LMMs’ self-attention mechanism. With the help of a CNN-based mask decoder and a SAM-based mask refiner, we achieve competitive visual grounding performance without losing any conversational ability of off-the-shelf LMMs. With both excellent conversational and visual grounding capabilities, the LMMs have the potential to perform complex visual perception and reasoning tasks like visual chain-of-thought reasoning.
6 Acknowledgements
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG3-PhD-2023-08-048T), the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).
Appendix A Appendix
In Sec A.1, we provide more detailed experimental results on both question-answering benchmarks and grounding benchmarks. Besides, we provide visualisation results in Sec A.3, including failure cases of existing grounding LMMs on general question-answering tasks, attention maps as well as segmentation masks, and examples of visual CoT and grounded conversations.
A.1 Benchmark results
Question-Answering Benchmarks. In addition to the four benchmarks reported in Table 1 of the main text, we also test the grounding LMMs on a wider range of question-answering benchmarks as shown in Table A1. Due to corrupted instruction-following abilities, existing grounding LMMs obtain zero or near-zero scores on these question-answering benchmarks.
Table A1: More evaluation results on question-answering benchmarks.
Model MME MMB MMVet LLaVAW POPE GQA VQAv2 AI2D
Existing Grounding LMMs
PixelLM-7B [58] 309/135 17.4 15.9 46.4 0.0 0.0 0.0 0.0
LISA-7B [28] 1/1 0.4 19.1 47.5 0.0 0.0 0.0 0.0
LLaVA-G-7B [77] - - - 55.8 - - - -
GLaMM-7B [57] 14/9 36.8 10.3 32.0 0.94 11.7 24.4 28.2
LaSagnA-7B [67] 0/0 0.0 16.7 34.5 0.0 0.0 0.0 0.0
General-Purpose LMMs
DeepseekVL-1.3B [42] 1307/225 64.6 34.8 51.1 88.3 59.3 76.2 51.5
MGM-2B [32] 1341/312 59.8 31.1 65.9 83.9 59.9 72.9 62.1
LLaVA-1.5-7B [37] 1511/348 64.3 30.5 69.0 85.9 62.0 76.6 54.8
HPT-Air-6B [62] 1010/ 258 69.8 31.3 59.2 87.8 56.2 74.3 64.8
HPT-Air-1.5-8B [62] 1476/308 75.2 36.3 62.1 90.1 59.4 78.3 69.0
MGM-7B [32] 1523/316 69.3 40.8 75.8 84.2 61.6 76.7 64.3
DeepseekVL-7B [42] 1468/298 73.2 41.5 77.8 88.0 61.3 78.6 65.3
LLaVA-1.6-7B [38] 1519/322 68.1 44.1 72.3 86.4 64.2 80.2 66.6
LLaVA-1.6-Mistral-7B [38] 1501/324 69.5 47.8 71.7 86.8 55.0 80.3 60.8
MGM-HD-7B [32] 1546/319 65.8 41.3 74.0 84.2 61.6 76.7 64.3
Referring Expression Segmentation. The results reported in Table 1 only include scores on the Val subsets of RefCOCO, RefCOCO+ and RefCOCOg. Here, we provide the grounding LMMs’ performances on all their subsets in Table A2. The metric used for Referring Expression Segmentation (RES) is cIoU.
Table A2: Detailed comparisons on Referring Expression Segmentation (RES).
Model RefCOCO RefCOCO+ RefCOCOg
val testA testB val testA testB val test
Specialised Segmentation Models
MCN [43] 62.4 64.2 59.7 50.6 55.0 44.7 49.2 49.4
LAVT [71] 72.7 75.8 68.8 62.1 68.4 55.1 61.2 62.1
GRES [36] 73.8 76.5 70.2 66.0 71.0 57.7 65.0 66.0
X-Decoder [82] - - - - - - 64.6 -
SEEM [83] - - - - - - 65.7 -
Existing Grounding LMMs
PixelLM-7B [58] 73.0 76.5 68.2 66.3 71.7 58.3 69.3 70.5
LISA-7B [28] 74.9 79.1 72.3 65.1 70.8 58.1 67.9 70.6
PerceptionGPT-7B [51] 75.1 78.6 71.7 68.5 73.9 61.3 70.3 71.7
LLaVA-G-7B [77] 77.1 - - 68.8 - - 71.5 -
GroundHog-7B [80] 78.5 79.9 75.7 70.5 75.0 64.9 74.1 74.6
GLaMM-7B [57] 78.6 81.1 76.1 70.5 74.9 63.0 74.8 74.8
LaSagnA-7B [67] 76.8 78.7 73.8 66.4 70.6 60.1 70.6 71.9
Grounding Frozen General-Purpose LMMs by F-LMM (Ours)
DeepseekVL-1.3B [42] 75.0 78.1 69.5 62.8 70.8 56.3 68.2 68.5
MGM-2B [32] 75.0 78.6 69.3 63.7 71.4 53.3 67.3 67.4
LLaVA-1.5-7B [37] 75.2 79.1 71.9 63.7 71.8 54.7 67.1 68.1
HPT-Air-6B [62] 74.3 79.4 71.8 64.0 71.7 57.2 67.5 68.3
HPT-Air-1.5-8B [62] 76.3 78.5 70.8 64.5 72.8 55.4 68.5 69.6
MGM-7B [32] 75.7 80.2 70.8 64.8 73.2 55.3 68.3 69.4
DeepseekVL-7B [42] 76.1 78.8 72.0 66.4 73.2 57.6 70.1 70.4
LLaVA-1.6-7B [38] 75.8 79.5 72.4 65.8 75.2 58.5 70.1 71.7
LLaVA-1.6-Mistral-7B [38] 75.7 79.6 71.2 66.5 75.5 58.1 70.1 70.3
MGM-HD-7B [32] 76.1 80.2 72.0 65.2 73.4 55.7 68.5 69.4
Panoptic Narrative Grounding. In Table 1 of the main text, we only report individual mask recalls on thing and stuff objects as well as the overall average recall. Here, we additionally report the mask recalls on singular and plural object nouns as shown in Table A3. As expected, segmenting plural nouns that refer to multiple object instances is more challenging for all the tested models.
Table A3: Detailed comparisons on Panoptic Narrative Grounding (PNG).
Model All Thing Stuff Singular Plural
Specialist Segmentation Models
MCN [43] 54.2 48.6 61.4 56.6 38.8
PNG [17] 55.4 56.2 54.3 56.2 48.8
PPMN [15] 59.4 57.2 62.5 60.0 54.0
XPNG [18] 63.3 61.1 66.2 64.0 56.4
Existing Grounding LMMs
PixelLM-7B [58] 43.1 41.0 47.9 49.1 27.7
GroundHog-7B [80] 66.8 65.0 69.4 70.4 57.7
GLaMM-7B [57] 55.8 52.9 62.3 59.7 45.7
Grounding Frozen General-Purpose LMMs by F-LMM (Ours)
DeepseekVL-1.3B [42] 64.9 63.4 68.3 68.3 56.1
MGM-2B [32] 65.6 64.4 68.4 69.1 56.9
LLaVA-1.5-7B [37] 64.8 63.4 68.2 68.2 56.1
HPT-Air-6B [62] 65.5 64.0 68.8 68.9 56.6
HPT-Air-1.5-8B [62] 65.4 64.1 68.5 68.9 56.5
MGM-7B [32] 66.3 65.3 68.6 69.8 57.3
DeepseekVL-7B [42] 65.7 64.5 68.5 69.2 56.7
LLaVA-1.6-7B [38] 66.3 65.1 69.0 69.8 57.3
LLaVA-1.6-Mistral-7B [38] 66.5 65.4 69.1 70.0 57.5
MGM-HD-7B [32] 66.7 65.6 69.1 70.1 57.8
A.2 Mask Decoder
Figure A1: The architecture of the mask decoder is based on a 3-stage U-Net [59] where the feature maps are downsampled and upsampled 3 times.
The architecture of the mask decoder based on a 3-stage U-Net [59] is shown in Figure A1, in which the feature maps are downsampled and upsampled three times. Downsampling encompasses two convolutional layers with a kernel size of 2 and 1, respectively. Upsampling is achieved using bilinear interpolation followed by two convolutional layers with a kernel size of 1. The number of parameters of the mask decoder is 8M.
A.3 Visualisation
Figure A2: Examples of grounding LMMs performing general question-answering tasks. The first example is obtained from MMBench [40], the second example is extracted from MME [16], and the last two examples are from LLaVA-in-the-Wild [39]. Existing grounding models (GLaMM, LISA, and PixelLM) fail to strictly follow user instructions nor correctly answer questions that necessitate a grasp of general world knowledge. In contrast, F-LMM (built upon LLaVA-1.5 [37] in the above examples), which completely inherits the conversational ability of general-purpose LMMs, performs excellently on these question-answering tasks.
General Multimodal Question-Answering. In Figure A2, we show some examples of grounding LMMs performing general question-answering tasks. When prompted to answer with single words (e.g., yes or no), existing grounding LMMs (GLaMM [57], LISA [28], and PixelLM [58]) usually fail to follow the user instructions. Besides, we also observe that the grounding LMMs tend to misunderstand the user’s questions as segmentation requests and reply mask tokens, e.g., ‘[SEG]’. Furthermore, these grounding LMMs fail to recognise the celebrity (Musk) and famous natural spot, exhibiting a worse grasp of world knowledge compared with a general-purpose LMM (e.g., LLaVA [39]). In contrast, F-LMM inherits the virtues of general-purpose LMMs (e.g., LLaVA-1.5 [37]) in instruction following and world knowledge comprehension, thanks to the ‘Frozen’ design philosophy.
Figure A3: Visualisations of K-Means results and segmentation masks. The attention weights (top-left) are mapped to 2D binary masks by the mask decoder (top-right) and then further optimised by the mask refiner (bottom-left). The samples are taken from the validation split of the PNG dataset [17]. The model used is DeepseekVL-1.3B [42].
From Attention Maps to Segmentation Masks. We visualise the geometric and spatial cues necessary for visual grounding in the word-image attention maps by applying KMeans clustering to the stacked attention maps that are collected from all transformer layers and attention heads. The attention maps of multiple-word objects are merged by element-wise average. As shown in Figure A3, we observe that the pixels of objects are roughly clustered together (top-left). With the CNN-Based mask decoder, the attention weights are mapped to 2D binary masks (top-right), which are then further optimised by the SAM-based mask refiner (bottom-left).
Figure A4: Visual Chain-of-Thought Reasoning. The model used is DeepseekVL-1.3B, and the samples are taken from the test set of VisCoT dataset [60]. The LMM is first prompted to think about the question-related object, which is then grounded by the mask head of F-LMM. The region of the question-related object is cropped and fed to the LMM to help answer the question.
Figure A5: Visualisations of grounded human-AI conversations. The key phrases or words in the conversations can be precisely localised by the mask head of F-LMM. The LMM used is DeepseekVL-1.3B.
Visual Chain-of-Thought Reasoning. Figure A4 shows more examples of visual CoT by F-LMM. The model used in these examples is DeepseekVL-1.3B. When the LMM is prompted to answer ‘What object is the most relevant to the question?’, the mask head of F-LMM grounds the LMM’s answer about the relevant object by generating a segmentation mask, the bounding box of which is used to crop the object region from the original image. Then, the cropped image region is fed to the LMM to obtain the final answer. As shown in Figure A4, the Visual CoT empowered by the LMM’s grounding ability is helpful when the LMM needs to focus on question-related regions for visual perception and reasoning.
Visualisation of Grounded Conversation. In Figure A5, we show some examples of grounding conversation by F-LMM. Our F-LMM maintains the LMMs’ original ability to follow the user’s instruction and understand unusual scenarios (e.g., the man ironing at the back of a taxi) while being able to localise the keywords and phrases during the conversations precisely. The model used in these examples is DeepseekVL-1.3B.
Appendix B Broader Impact
This paper addresses an important challenge in large multimodal models—improving the specialised performance while preserving the model’s general capabilities. By decoupling the grounding and conversational abilities, building upon the frozen LMMs, the proposed approach allows LMMs to visually ground objects and maintain their broad language capability. Our work is expected to have extensive benefits: (1) It enables the deployment of visually grounding LMMs in real-world applications that require both specialised multimodal capabilities and general language understanding, such as assistive tools and interactive robotics. (2) It paves the way for more flexible and adaptable multimodal AI systems that can be tailored to specific tasks or domains without compromising their core language capabilities. (3) Preserving instruction-following ability and resistance to hallucinations can improve the safety and reliability of the systems, making them suitable for high-stakes applications.
However, similar to many LMM-based systems, there are also potential negative impacts that should be considered: (1) Potential Bias: The pre-trained off-the-shelf LMMs used in the F-LMM approach may already contain biases, which could be propagated through the grounding process. (2) Potential for Displacement of Human Labor: The increased capabilities of visually grounding LMMs could lead to the displacement of human labor in certain domains, such as customer service, content creation, or image analysis. (3) Privacy and Ethical Concerns: Integrating visual grounding capabilities with language models raises privacy concerns, as the models could potentially be used to identify individuals or extract sensitive information from images.
To avoid misuse of the model, we will adopt the following safeguards: 1) Access Controls: Strict authentication and authorisation mechanisms will be implemented to ensure that only authorised and responsible individuals or organisations can access and use the models. 2) Usage Policies and Agreements: Clear usage policies and agreements will be established to define the intended purpose of the models. These policies will explicitly prohibit any malicious or harmful activities. Users will be required to agree to these policies and may face legal consequences if they violate them. 3) Transparency: We are committed to promoting transparency by providing comprehensive descriptions of the model’s capabilities, limitations, the training pipeline, and the datasets used.
Appendix C Limitations
While the proposed F-LMM approach demonstrates promising results in preserving conversational abilities while enhancing visual grounding, there are several key limitations that warrant consideration. (1) Inherited Biases and Limitations: As the F-LMM method is built upon frozen pre-trained LMMs, it inherits any biases or limitations present in the underlying models. These could include demographic biases, skewed knowledge representations, or other undesirable properties. (2) Limited Modality Scope: This work primarily focuses on vision-language multimodal interactions, without exploring other important modalities such as video, audio, and 3D point clouds. Expanding the scope to these additional modalities is a great direction to explore in the future. (3) Model Size Constraints: The experiments were restricted to LMMs up to 8 billion in parameter counts due to limited computing resources. Larger and more powerful models beyond this scale were not included. To address these limitations, future research could focus on mitigating biases, expanding the modality scope, and exploring larger-scale models.